OpenStudy (jiteshmeghwal9):

Given:-:-\[\alpha \beta \] are the roots of the equation \[x^2+px+q=0\] then the value of \[\left[ {\alpha \over \beta}+{\beta \over \alpha} \right]\] is equal to:-

OpenStudy (anonymous):

You can rewrite what you have to find: \[= \frac{\alpha^2 + \beta ^2}{\alpha. \beta}\]

OpenStudy (anonymous):

\[\frac{\alpha}{\beta}+\frac{\beta}{\alpha}=\frac{\alpha^2+\beta^2}{\alpha \beta}=\frac{(\alpha+\beta)^2-2 \alpha \beta}{\alpha \beta}\]

OpenStudy (jiteshmeghwal9):

Oh! I got it thanx a lot @waterineyes & @mukushla

OpenStudy (anonymous):

Welcome dear..

OpenStudy (jiteshmeghwal9):

The only thing is we have to put the values of alpha & beta in the equation, isn't it???

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

Sum of roots : \[\alpha + \beta = -p\] Product of roots: \[\alpha. \beta = q\] Just plug in the formula that mukushla gave you..

OpenStudy (jiteshmeghwal9):

k!